OpenStudy (anonymous):

use the distributive property to simplify 7 x17

OpenStudy (lncognlto):

Could you please clarify the question?

OpenStudy (anonymous):

the question just says Use the distributive property to simplify the expression 7 x 17

OpenStudy (asnaseer):

the distributive property states:\[a(b+c)=ab+ac\]use this to simplify the multiplication

OpenStudy (asnaseer):

think of how you can re-express 17 as the sum of two numbers

OpenStudy (anonymous):

I mean I tried doesnt it but it doesn't give me the right answer 7 x20= 140 7x7=49 140+49=189

OpenStudy (asnaseer):

first express 17 as the sum of two numbers - one of which should be a number that is very easy to multiply by

OpenStudy (anonymous):

the answer choices are 238, 97, 87, 119

OpenStudy (asnaseer):

what you did was 7*27 = 7(20+7) = 7*20 + 7*7

OpenStudy (anonymous):

I guess 20 because 17 is closer to 20

OpenStudy (asnaseer):

so - what two numbers add up to give 17?

OpenStudy (anonymous):

10+7

OpenStudy (anonymous):

ohhhhhhhhhhh

OpenStudy (anonymous):

I see

OpenStudy (asnaseer):

correct, so you can either say: 1) 17 = 10 + 7 or: 2) 17 = 20 - 3 which give you either: 1) 7 * 17 = 7(10 + 7) = 7*10 + 7*7 = ... 2) 7 * 17 = 7(20 - 3) = 7*20 - 7*3 = ...

OpenStudy (asnaseer):

yw :)